A man borrows $5,000 from the bank. Interest is charged at the end of each month 3% per month. He promises to repay at the end of two years the full amount plus interest in one big payment.
1. Calculate the total interest at the end of two years.
2. How much would be repaid at the end of two years?
3. If he decides to repay $150 per month, how much will he still owe the bank after the first 2 repayments?
4. Will he able to repay his debt if he continues to repay at $150 per month?

Answer by addingup(3677)

simple or compound interest?
Simple: Prt = interest, 5000(0.03*24) = 3600 this is the interest
5000 + 3600 = 8600 this is the amount to repay after 2 years
Monthly, his interest amounts to 5000*0.03 = 150 so if he pays 150/month he would only be paying the interest and would never get to repay the 5000